Grade 10General Physics

unit of power is 106 erg/minute, unit of force is 1000dyne,and the unit of time is 10-1 s.then the unit of mass in this system is

meenu s nair
8 Years agoGrade 10

1 Answer

Arun
8 Years ago
Dear student
 
Expression for power
P = F* v
P = F * L/t
Hence
L = P * t/F
 = 5/3 cm
Expression for force 
F = ma = m * v/t = m* L / t²
m = F * t² /L
 = 6 gm
 
Regards
Arun (askIITians forum expert)